OpenStudy (anonymous):

Age Problem: John is 2 times older than Jay. The difference of their ages is 12. How old is John??? What is the answer please????

OpenStudy (hoblos):

x= age of john y= age of jay x=2y x-y=12 replace equation 1 in equation 2 2y-y=12 y=12 then x=2y=24

OpenStudy (anonymous):

John - Jay = 12 2Jay - Jay = 12 -> Jay = 12 => John = 2 Jay = 2 * 12 = 24